### v1

# AI-Assisted Rapid Crystal Structure Generation Towards a Target Local Environment

Abstract
In the field of material design, traditional crystal structure prediction approaches require extensive structural sampling through computationally expensive energy minimization methods using either force fields or quantum mechanical simulations. While emerging artificial intelligence (AI) generative models have shown great promise in generating realistic crystal structures more rapidly, most existing models fail to account for the unique symmetries and periodicity of crystalline materials, and they are limited to handling structures with only a few tens of atoms per unit cell. Here, we present a symmetry-informed AI generative approach called Local Environment Geometry-Oriented Crystal Generator (LEGO-xtal) that overcomes these limitations. Our method generates initial structures using AI models trained on an augmented small dataset, and then optimizes them using machine learning structure descriptors rather than traditional energy-based optimization. We demonstrate the effectiveness of LEGO-xtal by expanding from 25 known low-energy sp2 carbon allotropes to over 1,700, all within 0.5 eV/atom of the ground-state energy of graphite. This framework offers a generalizable strategy for the targeted design of materials with modular building blocks, such as metal–organic frameworks and next-generation battery materials.
Keywords: Materials Informatics, Crystal Structure Prediction, Machine Learning
I INTRODUCTION
For new materials design, the ability to predict the crystal structure of a material is crucial for understanding its properties and potential applications. However, crystal structure prediction (CSP) has long been considered a grand challenge in materials science. Analogous to protein structure prediction Jumper et al. (2021), solving a typical CSP problem requires searching for the minimum energy arrangement of 10–100 atoms within an unknown periodic crystal lattice Oganov et al. (2019). This formulation results in a combinatorial search space of astronomical size Pickard and Needs (2011); Oganov and Glass (2006); Oganov, Lyakhov, and Valle (2011); Stillinger (1999); Martiniani et al. (2016). Additionally, evaluating each trial solution involves time-intensive quantum mechanical simulations, making the problem even more formidable.
Around the 2000s, the CSP field experienced a rapid development of non-empirical structure search algorithms by borrowing ideas from the global optimization community Oganov et al. (2019). Several algorithms—including random search Pickard and Needs (2011), genetic/evolutionary algorithms Oganov and Glass (2006); Wang et al. (2010); Lonie and Zurek (2011); Revard, Tipton, and Hennig (2018), simulated annealing Pannetier et al. (1990), and basin hopping Banerjee et al. (2021) were successfully applied to various material systems under both ambient and extreme conditions. These successes were largely attributed to strategies aimed at reducing dimensionality at the algorithmic level Pickard and Needs (2011); Oganov et al. (2019); Han et al. (2025). While the number of possible structural arrangements is infinite, many can be immediately discarded due to unfavorable energy costs. By incorporating reasonable initial guesses (e.g., imposing chemical and symmetry constraints Pickard and Needs (2011); Han et al. (2025)) and efficient structure-switching mechanisms (e.g., following physical vibration modes Lyakhov, Oganov, and Valle (2010); Zhu, Oganov, and Lyakhov (2012); Zhu et al. (2015), and self-learning intelligence Oganov and Glass (2006); Wang et al. (2010)), researchers were able to focus on exploring low-energy basins with a much reduced computational cost. However, these traditional approaches remain heavily reliant on high-end electronic structure prediction methods, which come with a computational cost scaling relation $O(N^{3})$ with respect to the number of valence electrons of the given system. As a result, a practical CSP task is limited to 20–30 atoms per periodic system, requiring several days to weeks on a single CPU node consisting of 48–96 cores.
Recently, the advent of artificial intelligence (AI) generative models has opened up new avenues for rapid crystal structure generation. Inspired by the successes of synthetic image generation and AI-driven protein structure prediction Jumper et al. (2021), various AI-based generative models have been introduced to the materials community since 2018, including Generative Adversarial Networks (GANs) Kim et al. (2020); Zhu et al. (2024), Variational Autoencoders (VAEs) Noh et al. (2019), and more recent active learning Merchant et al. (2023), diffusion Xie et al. (2021); Jiao et al. (2023); Zeni et al. (2025); Levy et al. (2024); Jiao et al. (2024); Miller et al. (2024); Zhong et al. (2025) and transformer models Antunes, Butler, and Grau-Crespo (2024); Gruver et al. (2024); Cao et al. (2024); Kazeev et al. (2025). These models offer a fundamentally new approach to exploring structure space by learning data-driven representations, allowing for the efficient generation of low-energy crystal structures at a much faster rate than traditional global optimization methods. However, the direct application of AI-based models to the CSP task faces a significant challenge: the generated structures often lack diversity or fail to satisfy essential physical constraints, as these models tend to either replicate structural prototypes from the training data, or generate invalid and unstable structure, rather than producing truly novel and physical-plausible structures Cheetham and Seshadri (2024). This limitation arises because these data-driven models primarily generate crystal structures only based on the statistical patterns from the training data source without incorporating the unique characteristics of crystalline materials, particularly their symmetries. To address the challenge, several recent works have attempted to incorporate symmetry into generative AI model design with the aim of enhancing the quality and validity of generated structures Jiao et al. (2024); Levy et al. (2024); Antunes, Butler, and Grau-Crespo (2024); Cao et al. (2024); Kazeev et al. (2025); Zhu et al. (2024). However, most of the previous works simply add each individual symmetry representation into the existing AI models. Consequently, the learned models may be able to generate new candidate structures with a symmetry identical to that in the training samples. Nevertheless, they are less likely to generate those closely symmetry-related structures due to the distinct symmetry representations, thus largely limiting the model’s learning efficiency.
In this work, we present a novel symmetry-informed generative model called LEGO-xtal (Local Environment Geometry-Oriented Crystal Generator) that overcomes the limitations of existing AI models in crystal structure prediction. Our approach is designed to generate realistic crystal structures with well-defined local environments (i.e. specified chemical building blocks) while preserving the inherent symmetries and periodicity characteristic of crystalline materials. As a proof of concept, we train our generative model on an augmented dataset of known three-coordinated carbon allotropes —commonly referred to as sp2 carbon structures— which enables the exploration of a broad range of structural configurations. Starting from only 25 known low-energy sp2 carbon allotropes (as well as additional 115 high-energy examples), we demonstrate the effectiveness of our approach by generating over 1,700 new structures with varied dimensionality, topology and unit cell sizes.
The remainder of this paper is organized as follows: First, we discuss the practical challenges in using AI models for crystal structure prediction and introduce our key solutions. Then, we present the detailed methodology of the LEGO-xtal framework. Finally, we analyze the generated structures and highlight several interesting findings.
II Computational Methodology
Unlike computer vision objects like images and point clouds that only require limited local information, crystal structures exhibit distinct long-range translational and local rotational symmetries. These symmetry elements are rigorously defined through crystallographic space groups using group theory Souvignier et al. (2016). When combined with specific chemical systems, crystal chemistry provides valuable rules and heuristics to guide our understanding of crystal packing and structure formation Oganov et al. (2019). To design an effective structure prediction model, we need to carefully consider how to incorporate these crystallographic and chemical principles into the modeling framework.
II.1 Symmetry-based Crystal Representation
[…]
Figure 1: Illustration of symmetry encoded crystal representation. (a) shows the example diamond crystal structure, (b) lists the associated Wyckoff symmetry operations to describe the carbon atoms in the diamond; (c) lists the complete tabular representation of a diamond crystal according to its highest space group symmetry 227.
Using this notation, it is straightforward to consider each structure in a tabular format with the following attributes.
  • Space Group number which ranges from 1 to 230.
  • Cell Parameters: $a,b,c,\alpha,\beta,\gamma$, which is a 6-vector that defines the cell lengths and angles in a standard crystallographic setting.
  • Wyckoff site 1: (index, element type, $x$, $y$, $z$), which represents the index of Wyckoff site and the associated fractional coordinates bounded in the (0, 1) interval.
  • Wyckoff site 2: (index, element type, $x$, $y$, $z$).
  • $\cdots$
Thus, one essentially needs two sets of variables to describe a structure: (i) the discrete variables denoting the exact space group symmetry and Wyckoff site indices; and (ii) the continuous variables describing the cell parameters and Wyckoff positions. For high symmetry cases, the cell parameters and $(x,y,z)$ coordinates may have additional constraints. For instance, cubic unit cells must have $a=b=c$ and $\alpha=\beta=\gamma=90^{\circ}$. The Wyckoff position may be completely or partially restricted to specific fractional coordinates due to site symmetry requirements. If input variables do not satisfy symmetry requirements, deterministic algorithms can check if variables can be symmetrized within an acceptable threshold. Considering the symmetry constraints, the free continuous variables are thus reduced to a 1-length vector [3.567].
Therefore, we can use a fixed-length 1D vector to represent a crystal structure with a maximum number of Wyckoff positions. For a maximum of 8 Wyckoff positions, the vector length is 1 + 6 + 8$\times$5 = 46. For a single-component system, this reduces to 39.
Although this representation is straightforward, it is not mathematically unique. In principle, one crystal structure can be represented by many different 1D vectors by choosing different choices of $(x,y,z)$ coordinates in a given Wyckoff site. To explicitly account for the symmetry, we can augment the representation by considering all possible multiplicities in each Wyckoff site as shown in Figure 1c.
II.2 Practical Challenges of AI Generative Models for Crystal Structure Prediction
Using this representation, the objective of crystal structure prediction is to find or generate the best combination of these variables that leads to a low-energy configuration. Inspired by the successes of AI generative models for image, point cloud, or text, we aim to develop a generative model to learn the statistical distribution from the training data for both discrete and continuous variables, and then generate new tabular data to represent novel crystal structures. However, there are several major challenges that prohibit the direct application of existing AI generative models to crystal structure generation.
First, there is a risk of lacking sufficient training data to describe the design space for the discrete variables (e.g., space group and Wyckoff site choices). In the context of crystal structure prediction, most materials generation tasks require the researchers to generate likely candidates from only a limited choices of good examples. However, with a total of 230 space groups and 1731 Wyckoff site choices, even for a simple unit cell of 10 atoms, the design space becomes astronomically large. With limited training data (typically ranging from several dozen to ten thousands examples), models are more likely to struggle to learn the underlying distribution of crystal structures. This lack of diversity and coverage can lead to overfitting and poor generalization, resulting in generated structures that are too similar to the training data or fail to explore the design space effectively.
Second, assuming the discrete design space (i.e. space group and Wyckoff site choices) has been sufficiently covered, the next challenge is to generate a valid crystal structure under the symmetry constraints. Currently, there exist two common practices to generate the continuous variables: (i) directly applying the diffusion model to generate the continuous variables for unit cell and Wyckoff sites Levy et al. (2024); Jiao et al. (2024); or (ii) combining random sampling with energy-based optimization to search for the low-energy structures Kazeev et al. (2025); Zhu et al. (2024). The first approach still relies on the availability of sufficient training data and may not be applied to small data tasks. The second approach is more reliable. However, for a specific space group and Wyckoff site choice, we may need to sample a large number of random configurations and then optimize them with the objective of finding the nearest lowest-energy configuration. This is often time-consuming and the quality of the optimized solution is overly reliant on the initial guess.
Figure 2: Proposed symmetry augmentation and pre-relaxation approaches to enhance the learning of crystal structure representation. (a) shows the process to augment the representation of diamond crystal by using its associated subgroup symmetry structures. (b) illustrates a graphic example to pre-relax the generated crystal to the desired sp2 local environment by using the power spectrum SO(3) descriptor.
II.3 Symmetry and Geometry Considerations
To effectively address these fundamental challenges in crystal structure prediction with AI generative models, we propose two key numerical recipes that significantly improve the model’s performance as illustrated in Figure 2.
[…]
Pre-relaxation with respect to the reference local environment. Once the model is capable of suggesting a promising candidate crystal, we then need to optimize the remaining continuous variables to get the realistic structure, under the conditioned space group symmetry constraint. In many practical applications, we may already know how the atoms may behave in a short-range cutoff. For instance, carbon may take sp, sp2, or sp3 local bonding; Porous materials such as zeolites and metal–organic frameworks (MOFs) are often categorized by their well-defined molecular building units, making them ideal targets for generative approaches that preserve local chemical environments. Therefore, we can utilize this knowledge to predefine the reference environment according to the recently developed power spectrum descriptors from spherical harmonics expansion Steinhardt, Nelson, and Ronchetti (1983); Kazhdan, Funkhouser, and Rusinkiewicz (2003); Bartók, Kondor, and Csányi (2013); Zagaceta, Yanxon, and Zhu (2020); Yanxon et al. (2020), and then relax the structure with respect to this reference environment. Figure 2b explains the idea of using SO(3) descriptor based reference environment for a randomly generated symmetric crystal. The numerical details can be found in the supplementary materials. Compared to the commonly used energy-based optimization, it not only saves a lot of computational time, but also effectively further reduces the searching space by excluding the possibility of handling low-energy configurations with an undesired local environment. Hence, we can rapidly screen if the trial structure can be relaxed to the desired environment prior to further consideration with more expensive energy-based optimization. Note that similar ideas of optimizing the structures at the descriptor space have also been explored in several recent works Tao, Shao, and Zhu (2024); Pickard (2025).
Using the aforementioned approaches, we expect that the search space can be greatly reduced for both discrete and continuous variables that are needed to describe the crystal. Therefore, one can manipulate the structural complexity with only a few variables, and the generative model can be trained on a small dataset with a limited number of known examples.
III The LEGO-xtal framework
Figure 3: The LEGO-xtal framework and its application to search for low-energy carbon allotropes. (a) shows the detailed pipeline of LEGO-cryst for adaptive generation of complex crystal structures. (b) summarizes the energy distribution of training dataset and newly generated structures by using this framework.
Motivated by preceding discussions, we propose the Local Environment Geometry-Oriented Crystal Generator (LEGO-xtal) framework that is capable of rapidly generating many feasible crystal structures from a few known examples. As illustrated in Figure 3, it consists of four main components: (1) data collection and augmentation, (2) generative modeling training and sampling, (3) structural pre-relaxation, and (4) energy ranking with physical models. Below, we will describe the details of each component in the context of generating complex structures comprising exclusively sp2-hybridized carbon atoms. Although the ground state of sp2-carbon is well known to be the graphite structure, many metastable carbon allotropes exist with intriguing properties. The exploration of these sp2 carbon allotropes has been highly sought after due to their fascinating electrical, optical, magnetic, thermal, and chemical properties Hoffmann et al. (2016); Lenosky et al. (1992); Pun and Miao (2018); Braun et al. (2018). Hence, we will use the sp2-carbon allotropes as a testbed to demonstrate the effectiveness of the LEGO-xtal framework.
III.1 Training Data Augmentation & Tabularization
First, we need to gather the already known data used for training as many as possible. For carbon research, the online Samara Carbon Allotrope Database (SACADA) database, which contains all available data about 3-periodic carbon allotropes extracted from scientific literature from Web of Science and Scopus databases Hoffmann et al. (2016). From SACADA, we extracted 140 carbon allotropes which are made of all sp2 carbon atoms. This includes a total of 126 unique (space group, Wyckoff sites) combinations, with energy spanning from 0 (graphite) to 1.344 eV/atom using the GGA-PBE pseudopotential Perdew, Burke, and Ernzerhof (1996) within the framework of density functional theory (DFT).
For each structure, we utilized PyXtal’s subgroup symmetry module Fredericks et al. (2021) to generate valid subgroup representations within constraints of 500 atoms and 8 Wyckoff sites per unit cell. Since the subgroup representation often results in larger unit cells and more Wyckoff sites, we excluded 36 structures with either low symmetry or large atom counts from the original 140 structures. For each of the remaining structures, we generated 2-15 subgroup symmetry representations. Additionally, we randomly sampled the multiplicity of each Wyckoff site due to symmetry operations. To evaluate the effectiveness of data augmentation, we created three datasets labeled as v1-60k, v2-120k, and v3-240k, with details summarized in Table 1.
Table 1: Summary of the training dataset for sp2 carbon allotropes.
Dataset   Number of samples Number of unique symmetries
Raw 140 127
v1-60k 63,114 419
v2-120k 122,322 449
v3-240k 242,766 485
In each representation, the space group number and Wyckoff site column values are always treated as discrete features. The cell parameters and Wyckoff site position values are treated as continuous features by default (cont-dataset). To improve the model’s learning efficacy, we also explore the possibility of transforming all continuous variables to discrete categorical values. In the discrete mode (dis-dataset), we divide the unit cell into a $100\times 100\times 100$ grid and convert each Wyckoff position into corresponding grid indices. For the cell parameters, we set up 100 bins between 0 and 35 for cell lengths, and 100 bins between 30 and 150 degrees for cell angles. The number of columns is 39 when the maximum number of Wyckoff sites is 8. Optionally, one may add some new columns, such as energy or other desired physical properties, to inject more information to guide the structure generation process.
III.2 Generative Model Training & Sampling
Prior to employing AI generative models, we transformed each row into an extended feature representation to enhance model learning. For discrete features, one-hot encoding is used, where each unique categorical value is mapped to an individual binary column. In this representation, a value of 1 indicates the presence of the corresponding category, while all other columns are set to 0. For continuous columns, we apply a Bayesian Gaussian Mixture Model (BGMM) approach as proposed by Patki et al Patki, Wedge, and Veeramachaneni (2016), in which each continuous value is first scaled to the interval $[-1,1]$, and then softly assigned to distinct Gaussian components via a probabilistic one-hot representation. This dual encoding strategy effectively captures both normalized numerical information and probabilistic cluster associations, improving the stability and efficacy of feature representation during generative model training. In the dis-dataset, we replace the BGMM-based encoding with fixed binning. All continuous features are discretized into categorical bins (e.g., cell lengths into 100 bins between 0–35 Å), and the resulting values are one-hot encoded. This yields a fully categorical representation for the entire dataset.
After transformation, the table expands to 525 columns in the cont-dataset and 2339 columns in the dis-dataset. These two encoding schemes allow us to benchmark model performance under both mixed-type and fully discrete data conditions. Next, we explored two generative models on both dis- and cont-type datasets to learn the underlying distribution of crystal symmetry data. The models are trained to synthesize new rows of tabular features that correspond to novel crystal structures that satisfy crystallographic constraints.
  • GAN. The Generative Adversarial Network (GAN) framework involves two neural networks—the Generator and the Discriminator—trained in a min-max game fashion, in which the Generator aims to create synthetic samples from random noise and fool the Discriminator, while the Discriminator attempts to distinguish real data from generated data. Our implementation is based on the Conditional Tabular GAN model Xu et al. (2019), which is specifically designed to handle tabular datasets with mixed data types (both continuous and categorical). During training, we adopt the Wasserstein GAN with Gradient Penalty framework Gulrajani et al. (2017) to improve training stability by enforcing the Lipschitz continuity condition via a gradient penalty. Once training is complete, the Discriminator is discarded, and only the trained Generator is used to produce new synthetic samples by using a random noise vectors sampled from a standard normal distribution.
  • VAE. The Variational Autoencoder (VAE) is a probabilistic generative model that learns to map input data into a structured latent space from which new samples can be generated Kingma, Welling et al. (2013); Rezende, Mohamed, and Wierstra (2014). The model consists of two neural networks: an Encoder that maps each input to the parameters (mean and variance) of a Gaussian distribution into latent space, and a Decoder that reconstructs the input from a sample drawn from this distribution. The latent space is regularized by enforcing the learned posterior distribution $q_{\phi}(z\mid x)$ to be close to a standard normal prior following the standard Normal distribution $\mathcal{N}(0,I)$, ensuring smoothness and continuity in the latent space which facilitates sampling and interpolation.
    The training objective includes a reconstruction loss and a Kullback-Leibler (KL) divergence term. The reconstruction loss ensures fidelity between the input and output (negative log-likelihood for continuous features, cross-entropy for categorical ones), while the KL term encourages the approximate posterior to match the prior, acting as a form of regularization. During the inference phase, new samples are generated by drawing random noise vectors from the standard multivariate normal distribution. These latent vectors are passed through the trained Decoder network, which maps them to the transformed tabular representation of crystal symmetry features.
For more details regarding GAN and VAE models, please refer to the Supplementary Materials Section 1 and Figs. S1 and S2). In a typical experiment, we train each GAN or VAE model with 250-1000 epochs, which is sufficient to converge the loss function values. Then we take the epoch that achieves the lowest loss function value to generate 100k - 500k samples for further consideration.
III.3 Structure Relaxation to Targeted Local Environment
Next, the structures yielded from the trained models need to be relaxed to the nearby energy minimum with the desired geometries. Instead of using the conventional structure optimization techniques based on classical force fields or electronic structure theory, we employed a descriptor-guided optimization framework that minimizes the dissimilarity in local atomic environments relative to a reference structure.
To represent the local atomic environment, the SO(3) descriptor $P$, with the option to include the explicit radial distribution function, is calculated to extract the radial and angular features of the neighbor density function Yanxon et al. (2020). For the purpose of this work, we selected a sp2 carbon (i.e., one centered atom is connected to 3 neighboring carbon with 1.42 Å  bond length and 120 bond angles in the same plane) to compute its power spectrum $P_{\text{ref.}}$ up to a radial cutoff of 2.0 Å. For each generated structure, we first evaluated the corresponding power spectrum $P$ for each atom, and then sought to minimize the summed mean squared error (MSE) loss function as follows:
$F_{\text{objective}}=\sum^{\text{all atoms}}_{i}\|P_{i}-P_{\text{ref.}}\|^{2},$ (1)
which quantifies the structural deviation from the standard sp2 environment. The goal of the optimization procedure is to adjust reduced variables (e.g., lattice parameters and free $x,y,z$ Wyckoff positions) to minimize this MSE loss function value, thereby steering structures toward graphite-like local order without requiring explicit energy evaluations. To perform the minimization, we utilized available optimization algorithms from the scipy.optimize.minimize library Virtanen et al. (2019), including Nelder-Mead Nelder and Mead (1965), L-BFGS-B Zhu et al. (1997), and Basin-Hopping Wales and Doye (1997).
During optimization, symmetry constraints were naturally enforced (due to the use of reduced variables) to ensure chemical realism and structural validity. Combining with the subgroup representation, this descriptor-guided strategy not only shortens the computational overhead associated with iterative energy evaluations but also allows the system to escape undesirable local minima associated with conventional force-based optimization. For more details, please refer to Fig. S3 and a motivating example in the Supplementary Section.
III.4 Energy Ranking & Database
After the pre-relaxation, the valid structures need to be evaluated by the energy ranking. Given that many structures needs to be processed and it is computationally expensive to perform quantum mechanical calculations for each structure, we considered two alternatives, (i) a reactive force field (ReaxFF) Van Duin et al. (2001) via the GULP package Gale (1997), a bond-order-based force field that can efficiently model chemical reactions and interatomic interactions; and (ii) a pretrained MACE model Batatia et al. (2022) via ASE Larsen et al. (2017), which utilizes higher order equivariant message passing for fast and accurate predictions of interatomic interactions in a large chemical space. Both methods are significantly faster than conventional DFT calculations, making them suitable for large-scale screening of generated structures.
Figure 4: Comparison of simulated energies for the selected 140 sp2 allotropes from the SACADA database.. (a) VASP vs. MACE and (b) VASP vs. ReaxFF. The dashed line indicates the ideal 1:1 correspondence.
Figure 4 compares energies calculated using MACE and ReaxFF models against reference DFT calculations performed with the VASP code Kresse and Furthmüller (1996) using the r2SCAN meta-GGA functional Furness et al. (2020a, b) and a KSPACING value of 0.15 Å-1. Generally, the MACE model shows a stronger correlation with DFT energies compared to ReaxFF, particularly for low-energy structures, making it slightly more suitable for our purposes. However, both models may stabilize structures into nonphysical configurations after pre-relaxation. Therefore, cross-checking generated structures with multiple models helps ensure physical validity. In our pipeline, we first use MACE for relaxation and energy ranking, then apply ReaxFF optimization to get the energy value as an alternative metric to infer the structural stability. According to our empirical observation, a feasible structure should be characterized by favorable energy ranking in both MACE and ReaxFF models.
The final database includes structures with MACE energies below 0.5-1.0 eV/atom. Beyond energy calculations, we characterize structural topology using CrystalNet.jl Zoubritzky and Coudert (2022) to identify novel configurations compared to existing databases using the common RCSR notations O’keeffe et al. (2008), providing additional insight into the structural diversity and stability of generated structures.
III.5 Approximated Time Costs
The rough timing information is shown in Table 2. Among all stages, the data collection and tabularization step can be done within several minutes on a single CPU. The time cost for model training is about 10-20 minutes using an H100 GPU for 250 epochs, and no more than 1 minute to generate 1k samples. The most significant time is spent on the pre-relaxation stage, as it needs to handle several hundred thousand samples with multiple iterations of optimization. On an AMD EPYC 9654P 96-Core Processor (each in 2.4 GHz), it takes 5-15 minutes per 1k samples and 8-16 hours for a total of 100k samples. Finally, the energy-based relaxation and ranking with MACE and ReaxFF require 5-20 minutes per 1k samples on the same 96-core processor, depending on the structure complexity and convergence behavior.
Table 2: Summary of time costs for each stage in LEGO-xtal.
Stage Device Speed
Data Augmentation Single CPU 3-5 mins
Training H100 GPU 10~20 mins / 250 epochs
Sampling H100 GPU 1~2 mins / 1k samples
Pre-relaxation AMD 96-core 5~15 mins / 1k samples
MACE-relaxation AMD 96-core 5~20 mins / 1k samples
ReaxFF-relaxation AMD 96-core 10~20 mins / 1k samples
IV Results & Discussions
To evaluate our framework, we trained multiple GAN and VAE models for generating sp2 carbon allotropes, by experimenting with various hyperparameters (e.g., training dataset size, continuous versus discrete representation, neural network architectures, number of training epochs). From each model, we generated 200k samples for evaluation. After pre-relaxation, we only retained structures exhibiting valid sp2 environments for further processing using MACE and ReaxFF. Below, we first analyze how different hyperparameter choices impact model performance, followed by a detailed examination of the sp2 allotrope database generated using LEGO-xtal.
IV.1 Model Performances Evaluation
First, Table 3 shows statistics of unique sp2 structures from different generative models trained on the v1-60k dataset with 200~500 epochs. Compared to the original 140 unique structures in the SACADA dataset, all models except GAN-Cont. successfully regenerated over 100 training structures within 100k samples, indicating a strong capability to learn from the training data. If we allow generating more samples, we expect all the training structures would be reproduced by these models. More encouragingly, each model generated a substantial number of unique structures (GAN-Cont.: 2937, GAN-Dis.: 4514, VAE-Cont.: 4862, VAE-Dis.: 4289). The limited overlap between models, as found in Table 3, suggests that each model learn different statistical patterns from the training data, leading to diverse generated structures.
Table 3: The overlap matrix showing the number of shared unique sp2 structures between 140 training structures and the 100k samples trained from v1-60k dataset using different models.
Train GAN-Cont. GAN-Dis. VAE-cont. VAE-Dis.
Train 140 62 111 100 132
GAN-Cont. 2937 649 848 707
GAN-Dis. 4514 979 952
VAE-Cont. 4862 1152
VAE-Dis. 4289
Table 4: Summary of metrics evaluated on 100k generated structures across different models with the best numerical results highlighted bold.
Model Dataset $N_{\text{valid\_xtal}}$ $N_{\text{valid\_env}}$ $N_{\text{unique}}$ $N_{\text{train}}$ $N_{\text{lowE\_all}}$ $N_{\text{lowE\_cubic}}$
GAN-Cont. v1-60k 95,313 10,687 2937 62 147 30
v2-120k 96,174 10,486 3132 65 40 31
v3-240k 96,404 9,621 3235 71 43 29
GAN-Dis. v1-60k 98,393 16,985 4514 111 212 62
v2-120k 98,861 18,406 4814 120 426 69
v3-240k 99,026 16,092 5,080 119 291 70
VAE-Cont. v1-60k 97,657 16,524 4,862 100 57 35
v2-120k 97,812 15,330 5,123 94 206 28
v3-240k 98,004 15,052 5,156 101 241 28
VAE-Dis. v1-60k 97,466 22,204 4,640 133 267 59
v2-120k 97,419 15,685 4,649 127 53 36
v3-240k 98,302 15,123 4,883 118 51 35
Next, we performed a more extensive benchmark to assess the quality and success of our generative method with different hyperparameters. To make a meaningful comparison between different models, we introduced the following metrics:
  • $N_{\text{valid\_xtal}}$: Number of generated samples that match the required constraints of the intended crystal symmetry.
  • $N_{\text{valid\_env}}$: Number of valid structures successfully optimized to adopt the sp2-type bonding configuration.
  • $N_{\text{unique}}$: Number of unique crystal configurations among the valid set, after removing duplicates.
  • $N_{\text{train}}$: Number of structures that were regenerated from the initial 140 training set.
  • $N_{\text{lowE\_all}}$ and $N_{\text{lowE\_cubic}}$: Number of unique (cubic) crystals with a MACE energy less than 0.55 eV/atom compared to the ground state, indicating thermodynamic stability.
These quantitative indicators collectively provide insights into how realistic, novel, and stable the generated structures are. The results, summarized in Table 4, show the relative performance of different models. Among them, the GAN models trained on continuous representation appear to perform notably worse than other models in nearly all metrics. This is likely due to the GAN’s difficulty in capturing the discrete nature of crystal symmetries and Wyckoff sites, which are crucial for generating valid crystal structures.
Nevertheless, other VAE and GAN models demonstrate similar performances across various metrics. On average, we observe that around 98% success rate in generating valid crystal structures that satisfy the space group and Wyckoff site symmetry constraints, and 15-22% of these structures can be relaxed to adopt the desired sp2 bonding configuration. The number of unique structures ranges from 3,235 to 5,156 per 100k samples, indicating a good diversity in the generated crystal configurations.
For realistic application, the more important metric is the number of low-energy unique structures. The GAN-Dis-v2/v3 model achieved higher numbers of low-energy (cubic) structures as compared to the corresponding v1 model, suggesting that more explicit subgroup symmetries can promote the model’s learning efficacy. A similar trend is also observed for the VAE-continuous models. However, this is not the case for the VAE-Discrete models, in which the model trained on the v1-60k dataset produced the highest number of low-energy structures. This may be due to the fact that the VAE-Dis. model uses the noise sampled from continuous Gaussian distribution, but the generated results should be represented in discrete values. Therefore, the VAE models may have some intrinsic limitations in learning the discrete variables.
While it is overall difficult to make a definitive recommendation on a single best model for this task, we observe a general trend that most models can rapidly generate a large number of low-energy sp2 carbon allotropes with diverse structures, as compared to the original 140 training structures. This opens up new opportunities for effectively harnessing the power of AI generative models to explore the structural space of complicated materials. The improvements of model performance will be left for the future, as it is beyond the scope of this work.
IV.2 The sp2 Carbon Allotrope Database
Figure 5: Complexity analysis of the 1741 low-energy sp2 carbon allotropes. (a)-(c) shows the distributions of the energy of the generated structures by number of atoms in the unit cell, space group symmetry, and number of reduced variables, respectively. (d) shows two representative low-energy structures demonstrating the intrinsic crystal searching challenge. In (d), the structures are labeled by their Pearson symbol and relative DFT energies as compared to the ground state graphite structure. The Pearson symbol is a shorthand notation that describes the crystal structure, including the space group and the number of atoms in the unit cell.
Since no single model significantly outperformed the others, we ran multiple models to comprehensively explore the structural space of sp2 allotropes. After removing duplicate structures, we identified a total of 21,786 structures within a 1.0 eV MACE energy window, including 1,741 structures with energies less than 0.5 eV above the ground state. In the original 140 data used in training, there exist 25 sp2 carbon allotropes that have the MACE energy less than 0.5 eV/atom higher than the ground state graphite structure (see Figure 3b). Using the LEGO-xtal, we have significantly expanded it to 1700+ distinct structures, evidently demonstrating the power of our LEGO-xtal framework to rapidly generate novel structures. To double check the accuracy of MACE energy ranking, we also manually selected 150 structures for further relaxation with DFT-r2SCAN using the VASP package. The complete database of 1,700+ low-energy sp2 carbon allotropes, as well as ReaxFF, MACE and DFT-r2SCAN energies, is available at https://lego-crystal.onrender.com. Below, we will focus on the chemical analysis of these low-energy structures.
IV.2.1 Distribution of low-energy crystals by complexity
[…]
In addition, it is important to emphasize that our approach does not necessarily only favor high-symmetry structures. As shown in Figure 5b, there exist many structures with lower symmetry space groups other than the cubic symmetry. Using the number of reduced crystal variables as a metric to probe the model’s predictive capabilities (Figure 5c), we find that the majority of low-energy structures have fewer than 20 reduced variables. This indicates that our framework is most effective at exploring the structural space of sp2 carbon allotropes with a limited number of variables, less than 20. However, we can still find that a notable portion of structures (280 out of 1741) have more than 20 reduced variables. Figure 5d highlights a representative low-energy structure with 23 reduced variables - a 3D-periodic structure containing 72 atoms per unit cell in $P\bar{3}m1$ symmetry. This finding demonstrates that our framework can effectively explore more complex structural spaces despite increasing uncertainty at higher dimensions.
Figure 6: The screened low-dimensional sp2 carbon allotropes.. (a) Distribution of the energy of the generated structures grouped by dimensionality. (b-d) show three representative 0D, 1D, and 2D structures labeled by the Pearson symbol and the relative DFT-r2SCAN energies as compared to the ground state graphite structure.
IV.2.2 The 0-2D low-energy sp2 carbon allotropes
Given the nature of sp2 bonding, layered graphite is unquestionably the ground state. Figure 6a shows the breakdown of energy distribution by detected dimensionality. The majority of low-energy structures (< 0.15 eV/atom) are characterized by stacked hexagonal graphite layers. In addition to the commonly known stacking polytypism due to layer translations Zhu et al. (2015), we also discovered many variations arising from layer twisting. Among these, the most interesting structure is shown in Figure 6d, which is an AB-type stacked layered structure with about a 30-degree rotation angle between two adjacent layers. This structure is not only low in energy but also exhibits a unique 2D periodicity. This finding suggests that our framework can effectively explore the structural space of sp2 carbon allotropes that is related to recent work on the twisted bilayer graphene structures with varying rotation angles Bistritzer and MacDonald (2011); Cao et al. (2018).
Beyond the stacking graphite-type layers, we also identified several other planar configurations in different types of ring topology (e.g., mixed five-, seven, eight-membered rings, see Fig. S4 in the supplementary materials) with an energy spanning from 0 to 0.5 eV range. Additionally, several unique 1D carbon nanotube-like structures with varied diameters (e.g., a representative model 1D-hR108 in Figure 6c) have been observed in our results. Finally, we also found two large 0D allotropes (see 0D-cP176 in Figure 6d) featured by the combination of five, six and eight-membered rings that are similar to the well known C60 and C70 Haymet (1985). Notably, these 0D allotropes are calculated to possess lower energies (0.2-0.3 eV/atom) than C60 and C70. Hence, they have better stability if they can be synthesized in the experiment. These examples demonstrate our framework’s capability to explore diverse structural motifs across multiple dimensionalities.
Figure 7: A list of low-energy negative curved graphite structures found by LEGO-xtal. The structures are labeled by the topology according to RCSR notation O’keeffe et al. (2008) (when available), the Pearson symbol, and relative DFT-r2SCAN energies as compared to the ground state graphite structure. The previously reported structures (a-d) are also marked by the asterisks.
IV.2.3 The 3D Negative Curved Graphite allotropes
Apart from the 300+ low-dimensional crystal packing motifs, the majority of low-energy structures possess extended networks with 3D periodicity. Not surprisingly, the lowest energy 3D sp2 allotropes closely resemble graphite-like layered structures (see Fig. S5 in the supplementary materials and previous literature Zhao, Zhang, and Guo (2020)). However, the conceptually more compelling 3D sp2 allotropes are those containing negative Gaussian curvature that form inward curved surfaces, resembling saddle shapes or hyperbolic geometry Pun and Miao (2018). Unlike C60 with a positive curved graphite network, these 3D negative curved graphite (NCG) structures are more complex and feature fascinating properties due to their unique topology. Since 1990, only 15 NCG structures have been proposed through either mathematical derivation or sophisticated physical modeling approaches Mackay and Terrones (1991); O’Keeffe, Adams, and Sankey (1992); Lenosky et al. (1992); Mackay and Terrones (1993); Townsend et al. (1992); Phillips et al. (1992); Tagami et al. (2014); Braun et al. (2018).
Within the 1741 structures, we have observed a total of 386 sp2 cubic structures. To confirm their stability, we selected the 100 NCG candidates for further energy ranking using VASP at the DFT level. And the top 8 ranked structures are listed in Figure 7. Among them, we found four new structures (NCG1-cP192, NCG2-cP108, NCG3-cP192, and NCG4-cI144) that have neither been used in training nor reported in the previous literature. Remarkably, both NCG1-cP192 and NCG2-cP108 are $\sim$0.015 eV/atom lower in energy than the previously known lowest-energy mct-cI192 NCG structure Phillips et al. (1992). Furthermore, we computed the phonon dispersions for the NCG1-NCG3 structures and confirmed their dynamical stability (see Fig. S6 in the Supplementary Materials).
In addition, there exist 41 structures with more than 300 atoms in the unit cell, highlighting that our approach can effectively handle complex structures that were seldom reported in the literature. Interestingly, by comparing with the structures used in training, we found that all newly identified structures exhibit some similar features to the training data while still maintaining unique characteristics. This indicates that our framework can effectively explore and expand upon known structural motifs rather than simply replicating existing structures. These remarkable findings demonstrate the power of our framework to discover new materials with complex geometries.
V Conclusions
In this work, we present LEGO-xtal, a novel framework that combines subgroup symmetry augmentation, descriptor-guided pre-relaxation, and advanced generative models to efficiently explore crystal structure space from limited training data. Starting from just 25 (140) known sp2 carbon structures, our approach successfully generated over 1,700 (21,786) diverse low-energy allotropes, including high-symmetry structures with large unit cells (up to 960 atoms) and unique 0D, 1D, and 2D configurations. This demonstrates the framework’s capability to discover complex crystal structures while maintaining physically realistic local environments.
The success of our framework stems from the synergistic combination of subgroup symmetry augmentation and descriptor-guided pre-relaxation. While symmetry augmentation enables efficient exploration of the structural space through reduced discrete variables, pre-relaxation ensures the generated structures maintain desired local environments at minimal computational cost. Together, these strategies dramatically reduce the dimensionality of the crystal structure search space.
Looking ahead, we aim to extend subgroup symmetry augmentation beyond single components and reference environments, enabling the generation of more complex structures with multiple components and varied local environments in a single model. This extension requires adapting pre-relaxation strategies to handle multiple reference environments, either by modifying the environmental loss function within traditional optimization routines Pickard (2025) or by employing more flexible diffusion models that incorporate local environment constraints during the training stage. These improvements will enable the generation of increasingly complex structures with higher dimensionality and more sophisticated local environments, which will be the focus of our future work.
Acknowledgments
This research was sponsored by the U.S. Department of Energy, Office of Science, Office of Basic Energy Sciences, and the Established Program to Stimulate Competitive Research (EPSCoR) under the DOE Early Career Award No. DE-SC0024866, the UNC Charlotte’s seed grant for data science, as well as European Union (ERDF), Région Nouvelle Aquitaine, Poitiers Univeristy, and French government programs “Investissements d’Avenir" (EUR INTREE, reference ANR-18-EURE-0010) and PRC ANR MagDesign and TcPredictor. The computing resources are provided by ACCESS (TG-DMR180040) and High-Performance Computing Centre Adastra/CINES of GENCI (projects A0140807539 and A0160815101). During the preparation of this work, the authors used GitHub Copilot in order to improve the code readability and documentation. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Data availability
The LEGO_xtal source code, instructions, as well as scripts used to calculate the results of this study, are available in https://github.com/MaterSim/LEGO-xtal. The complete list of 1741 low-energy sp2 carbon allotropes, as well as ReaxFF, MACE and DFT-r2SCAN energies, is listed in http://lego-crystal.onrender.com.
Conflict of interest
All authors declare that they have no conflict of interest.
Author contributions
The concept of Crystal Structure Generator From Chemical Building Blocks was initially discussed by G.F. and Q.Z. (with the latter being a Visiting Professor at Poitiers University in June 2023). Q.Z. and G.F. co-conceived the idea. With the help of G.F, Q.Z. initiated the framework. Q.Z., D.D., H.X., and G.F. supervised this project. O.G.R., M.S.R., H.X., and Q.Z. implemented the code, S.P. performed the electronic structure calculations. All authors analyzed the results and contributed to manuscript writing.